A calorimeter contains 35.0 mL of water at 14.5 ∘C . When 1.80 g of X (a substance with a molar mass of 62.0 g/mol ) is added, it dissolves via the reaction

X(s)+H2O(l)→X(aq)

and the temperature of the solution increases to 28.5 ∘C .

Calculate the enthalpy change, ΔH, for this reaction per mole of X.

Assume that the specific heat of the resulting solution is equal to that of water [4.18 J/(g⋅∘C)], that density of water is 1.00 g/mL, and that no heat is lost to the calorimeter itself, nor to the surroundings.
Answer #1

moles of X = 1.80 / 62 = 0.0290

total mass of solution = 35 + 1.80 = 36.80 g

Q   = m Cp dT

Q = 36.80 x 4.184 x 28.5

Q = 4388 J

Q   = 4.388 kJ

ΔH   = - Q / n

         = -4.388 / 0.0290

         = -151 kJ / mole

ΔH = - 151 kJ / mole